OpenStudy (anonymous):

The area A of a trapezoid is given by the formula A = 0.5(b + B)h, where b and B are the sizes of the bases and h the size of the height of the trapezoid. Express B in terms of A, b and h.

OpenStudy (nurali):

A = (1/2) (b + B) h 2A = (b + B)h 2A/h = b + B B = 2A/h - b

OpenStudy (anonymous):

how did you get (1/2) ?

OpenStudy (nurali):

1/2 means 0.5

OpenStudy (anonymous):

I understand, thanks ^^

OpenStudy (nurali):

Anytime.